the perimeter of a rectangle is 30cm. if the width is 5cm, what is it's length
Found 3 solutions by Theo, MathTherapy, Alan3354:
Answer by Theo(13342)

perimeter = 2 * length + 2 * width.

perimeter = 30 cm
width = 5 cm

equation becomes 30 = 2 * length + 2 * 5

simplify to get 30 = 2 * length + 10

subtract 10 from both sides of the equation to get:

30 - 10 = 2 * length

combine like terms to get:

20 = 2 * length

divide both sides of the equation by 2 to get:

20 / 2 = length

simplify and solve for length to get:

length = 10.

length = 10 and width = 5

2 * length + 2 * width = 2 * (length + width) = 2 * (15) = 30

solution is:

length = 10 cm.


Answer by MathTherapy(10552)

the perimeter of a rectangle is 30cm. if the width is 5cm, what is it's its length
With length being L and width, W, we get: 2L + 2W = 30_____L + W = 15
L + 5 = 15 ----- Substituting 5 for W

Answer by Alan3354(69443)

it's = it is